HIGH-RESOLUTION INFRARED SPECTROSCOPY OF JET-COOLED VINYL RADICAL: SYMMETRIC CH_{2} STRETCHING VIBRATIONS

\underline{FENG DONG}, and DAVID J. NESBITT, JILA, University of Colorado, Boulder, CO 80309.

Vinyl radical, C_{2}H_{3}, is a transient but fundamental hydrocarbon species which plays important roles in a number of combustion processes. Previous high resolution spectroscopic studies of this species have been very limited; only the CH_{2} wagging mode in the electronic ground state has been ^{a} studied with IR diode laser absorption methods. Guided by both recent time-resolved FTIR emission ^{b} and high level density functional predictions, we have recently used the slit jet discharge IR laser spectrometer to observe and assign high-resolution infrared spectra of jet-cooled vinyl radical in the symmetric CH_{2} stretch region. In addition to structural information, the tunneling dynamics due to in-plane rocking of the α-C-H will be discussed.

aH. Kanamori, Y. Endo, and E. Hirota, J. Chem. Phys. 92 (1), 197 (1990).

bL. Letendre, D. K. Liu, C. D. Pibel, J. B. Halpern, and H. L. Dai, J. Chem. Phys. 112 (21), 9209 (2000).